OpenStudy (anonymous):

what is the value of a pythagorean theorem of x-10 x+10

OpenStudy (jamesj):

So if the length of the perpendicular sides are a and b, what is the length of the hypothenuse? It's\[\sqrt{a^2 + b^2}\] So here a = x - 10 and b = x + 10. Substitute and evaluate.